# METANOIA: A Lifelong Intrusion Detection and Investigation System for Mitigating Concept Drift

### v1

@@ VI-A Experiment Protocol
We implemented a METANOIA prototype in Python. We utilized PyG [57] to generate reconstruction losses, Networkx [58] to construct the minimum graph, and GraphViz [59] to visualize the reconstructed attack scenario graph. METANOIA leverages the agents to collect data from monitored hosts. The collected data have the same format as other detection systems [15, 14, 12, 1].
We selected three representative anomaly-based APT online detection systems, namely UNICORN [1], PROGRAPHER [14], and KAIROS [15], as baselines for comparison. And we will compare the window-level detection accuracy with these three and node-level detection accuracy with the latter two. It is noteworthy that PROGRAPHER and KAIROS did not directly evaluate their node-level detection capabilities in the original paper. However, we observed that both reconstruct attack scenarios, aligning with METANOIA’s functionality. Therefore, we considered the nodes involved in reconstructed attack scenarios as the basis for evaluating node-level detection accuracy. We directly used the source code released by UNICORN and KAIROS and employed their default parameters to avoid possible biases. As for PROGRAPHER, since it is not open source, we reproduce it and will release its code. Our experiments were carried out on a server running Ubuntu 18.04 64-bit OS with Intel(R) Xeon(R) Silver 4214R CPU @ 2.40GHz, 256GB memory.
@@ VI-C Datasets
The purpose of METANOIA is to help long-running online detection systems combat the problem of false positives caused by concept drift that requires long-term datasets for evaluation. Given that there are multiple adversarial scenarios designed to mimic real-world Advanced Persistent Threats (APTs) on enterprise networks within the Transparent Computing (TC) program, we choose to employ datasets from DARPA’s TC. Specifically, multiple red teams (e.g., CADETS, ClearScope, and THEIA) launched a series of attacks toward critical services such as web, email, and SSH servers, while engaging in benign activities such as browsing websites, checking emails, and SSH log-ins. The datasets from the third (E3) and fifth (E5) engagements of DARPA TC are publicly accessible [60, 32].
Table III details the statistics from the different datasets, including duration time (specific dates and hours), the number of nodes and edges, the specific attack time and their proportions, as well as the number of attack-related nodes. The open-source report for DARPA TC E3 [60] and E5 [32] show an attack for E3-CLEARSCOPE on April 12, 2018 from 15:19 to 15:24, an attack for E5-CLEARSCOPE on May 17, 2019 from 15:42 to 16:00. However, we couldn’t find any relevant attack records (including file downloading and process creating) in the corresponding log files for this period. Therefore, we do not label it as attack time.
@@ VI-D RQ 1: The detection performance of METANOIA
KAIROS [15] is a window-level detection system, THREATRACE [13] is a node-level detection system, while UNICORN [1] and PROGRAPHER [14] are graph-level detection systems. To ensure fairness, we will compare METANOIA with three baselines at their respective levels to demonstrate the detection performance of METANOIA.
@@ VI-D2 Graph-level
UNICORN transforms streaming data into fixed-size, updatable graph sketches that preserve Jaccard similarity [61] to model a system’s behavior. Each sketch represents a snapshot, depicting an overall summary of the system’s state from the beginning of execution to the time of snapshot capture. PROGRAPHER [14], on the other hand, converts snapshots into fixed-size vectors using graph neural networks graph2vec [62] and leverages recurrent neural networks TextRCNN [63] to learn benign system behavior changes.
@@
For comparison, UNICORN and PROGRAPHER all perform worse than METANOIA in nearly every metric. We learned that there are two reasons why UNICORN’s performance is worse than reported in the original paper: (1) UNICORN was trained using benign datasets under DARPA TC that were not publicly available, leading to bias in the model. (2) UNICORN did not enforce that the testing graphs occur after training and validation, which is a typical Data Snooping [64]. While PROGRAPHER was not open-sourced, we replicated it based on the description of the paper. Keeping the hyper-parameters and necessary components unchanged, we believe the main reason for PROGRAPHER’s poorer performance compared to the original paper is that the encoder (Graph2Vec) cannot generate suitable embedding vectors. We primarily used Doc2Vec [65] provided by Gensim [66] in combination with RSG mentioned in PROGRAPHER to generate embedding vectors for snapshots, but the performance was unsatisfactory. We attribute the result to our lack of expertise rather than the original authors, and we intend to open-source the reproduction code of PROGRAPHER.
@@ VI-D3 Node-level
THREATRACE is a detection system that assesses the anomalousness of nodes using a multi-model framework. While METANOIA generates a reconstruction graph of the attack scenario corresponding to the malicious windows, allowing for evaluation at the node level for both. THREATRACE [13] is an open-source system, enabling us to run it directly for comparison. To maintain fairness, we adopted THREATRACE’s way of computing metrics and selected a subset of the DARPA TC dataset used by THREATRACE.
@@ VI-E RQ2: The Importance of METANOIA’s Components
Can the RN Pool be contaminated? From Table VII, we can observe that the constructed RN Pool has no intersection with the attack nodes across almost all datasets. This is attributed to METANOIA’s stringent suspicious state transfer rules in constructing the RN Pool, thereby enabling METANOIA to avoid learning malicious behaviors and introducing false negatives. Prior work [28, 9] suggests that suspicious state transfer captures the essence of intrusion attacks, and METANOIA utilizes this rule to inversely identify rehearsal nodes that can provide replay information of critical scenarios. Additionally, in the case of E3-ClearScope, due to the same reason (absence of events related to process creation), the suspicious state can’t transfer effectively, resulting in all attack-related nodes being included in the RN Pool. We attribute this to issues with the auditing system.
@@ VI-F RQ3: The Effectiveness of Reconstructing Attack Scenarios
Previous researches [15, 33, 67] indicate that reconstructing attack scenarios is an essential capability for existing detection systems. Unlike rule-based detection methods, anomaly-based detection systems focus more on this ability. The reason for this is that anomaly-based detection systems inherently have a certain amount of false positives, requiring involvement from security analysts. Reconstructing a complete and concise attack scenario can effectively assist security analysts in eliminating false positives and reducing time overhead.
@@ VII Discussion
Evasion Attack. Evasion attack refers to the actions taken by the attacker to mimic benign behavior to deceive detection systems during an attack. While evasion attacks pose a threat to all PIDSes, evading anomaly-based PIDSes is not straightforward. Therefore, to mimic benign behavior, attackers need to carefully orchestrate their attack actions, ensuring that malicious behavior exhibits benign features at the log level while maintaining their attack objectives. This requires attackers to have a deep understanding of both the target system’s benign behavior and the mechanics of running PIDSes. Previous work [17] has demonstrated that existing graph-based adversarial attacks cannot evade the detection of PIDSes because the provenance graphs have more structural features and time constraints. A recent robustness study [68] has indicated that evasion attacks can be achieved by automatically modifying system logs to evade existing PIDSes and have released evasion datasets based on DARPA E3-THEIA. However, KAIROS pointed out that this dataset only includes a small portion of the attack activities described in the DARPA ground truth, rather than the complete attack traces included in the original DARPA dataset. Further investigation confirmed this, and we did not proceed with further evaluation.
Poisoning Attack. If an attacker can poison the training data to contain malicious activity for model learning, then future attacks will go undetected. To the best of our knowledge, SIGL [17] is the only PIDS that evaluates its robustness against data poisoning in-depth, but only for the software installation scenario. ShadeWatcher [12] inserts attack data from one day during training and demonstrates that it can still detect the corresponding attack. We performed a similar evaluation and also obtained good detection results. METANOIA adopts a new anomaly detection paradigm that requires replaying scenarios to combat catastrophic forgetting and is therefore more sensitive to poisoning attacks. METANOIA starts from the opposite side of the suspicious information flow to avoid the discrimination paradox and thus combat the poisoning attack.
@@ VIII Related Work
In terms of detection methods, existing PIDSes can be categorized as heuristic-based and anomaly-based. Heuristic-based PIDSes [7, 69, 2, 8, 28] use empirical knowledge to construct matching rules to find known attacks in the provenance graph. However, they have difficulties in detecting unknown attacks and extending scale. Anomaly-based PIDSes detect intrusions by identifying deviations from normal behavior. Besides, offline PIDSes [12, 18, 17], due to their high operational overhead, fail to achieve efficient real-time intrusion detection. Online PIDSes [1, 13, 14] relies on a fixed-size training set and cannot effectively address the problem of concept drift. UNICORN [1] combats concept drift by gradual forgetting but causes key information to decay with duration. KAIROS [15] relies on the intervention of analysts to re-train the model to combat concept drift, resulting in breaking attack-agnostic and high training costs. Although KAIROS acknowledges this issue in the paper, it does not provide a targeted solution, leaving it for future work.
Moreover, previous PIDSes [30, 11, 10, 1] required manual investigation of large-scale provenance graphs, rendering them impractical for real-world use. Recently, THREATRACE [13] and SHADEWATCHER [12] began identifying anomalies at the node level, but they cannot reconstruct complete attack scenarios like METANOIA. SIGL [17] attempts to generate graphs for software installation, but it struggles to scale to full-system intrusion detection. PROGRAPHER [14], on the other hand, sorts the root subgraphs (RSGs) in snapshots by anomaly severity to help narrow down the scope of the investigation but loses the connection between multiple snapshots. Meanwhile, KAIROS [15] attempts to reconstruct attack scenarios in anomaly queues using community discovery algorithms but can only generate multiple star-shaped small attack scenarios (according to its open-source results).

### v2

@@ VI-A Experiment Protocol
We implemented a METANOIA prototype in Python. We utilized PyG [pyg] to generate reconstruction losses, Networkx [networkx] to construct the minimum graph, and GraphViz [graphviz] to visualize the reconstructed attack scenario graph. METANOIA leverages the agents to collect data from monitored hosts. The collected data have the same format as other detection systems [kairos, prographer, shadewatcher, han2020unicorn].
We selected three representative anomaly-based APT online detection systems, namely UNICORN [han2020unicorn], PROGRAPHER [prographer], and KAIROS [kairos], as baselines for comparison. And we will compare the window-level detection accuracy with these three and node-level detection accuracy with the latter two. It is noteworthy that PROGRAPHER and KAIROS did not directly evaluate their node-level detection capabilities in the original paper. However, we observed that both reconstruct attack scenarios, aligning with METANOIA’s functionality. Therefore, we considered the nodes involved in reconstructed attack scenarios as the basis for evaluating node-level detection accuracy. We directly used the source code released by UNICORN and KAIROS and employed their default parameters to avoid possible biases. As for PROGRAPHER, since it is not open source, we reproduce it and will release its code. Our experiments were carried out on a server running Ubuntu 18.04 64-bit OS with Intel(R) Xeon(R) Silver 4214R CPU @ 2.40GHz, 256GB memory.
@@ VI-C Datasets
The purpose of METANOIA is to help long-running online detection systems combat the problem of false positives caused by concept drift that requires long-term datasets for evaluation. Given that there are multiple adversarial scenarios designed to mimic real-world Advanced Persistent Threats (APTs) on enterprise networks within the Transparent Computing (TC) program, we choose to employ datasets from DARPA’s TC. Specifically, multiple red teams (e.g., CADETS, ClearScope, and THEIA) launched a series of attacks toward critical services such as web, email, and SSH servers, while engaging in benign activities such as browsing websites, checking emails, and SSH log-ins. The datasets from the third (E3) and fifth (E5) engagements of DARPA TC are publicly accessible [darpa3, darpa5].
Table III details the statistics from the different datasets, including duration time (specific dates and hours), the number of nodes and edges, the specific attack time and their proportions, as well as the number of attack-related nodes. The open-source report for DARPA TC E3 [darpa3] and E5 [darpa5] show an attack for E3-CLEARSCOPE on April 12, 2018 from 15:19 to 15:24, an attack for E5-CLEARSCOPE on May 17, 2019 from 15:42 to 16:00. However, we couldn’t find any relevant attack records (including file downloading and process creating) in the corresponding log files for this period. Therefore, we do not label it as attack time.
@@ VI-D RQ 1: The detection performance of METANOIA
KAIROS [kairos] is a window-level detection system, THREATRACE [threatrace] is a node-level detection system, while UNICORN [han2020unicorn] and PROGRAPHER [prographer] are graph-level detection systems. To ensure fairness, we will compare METANOIA with three baselines at their respective levels to demonstrate the detection performance of METANOIA.
@@ VI-D2 Graph-level
UNICORN transforms streaming data into fixed-size, updatable graph sketches that preserve Jaccard similarity [yang2017histosketch] to model a system’s behavior. Each sketch represents a snapshot, depicting an overall summary of the system’s state from the beginning of execution to the time of snapshot capture. PROGRAPHER [prographer], on the other hand, converts snapshots into fixed-size vectors using graph neural networks graph2vec [narayanan2017graph2vec] and leverages recurrent neural networks TextRCNN [lai2015textrcnn] to learn benign system behavior changes.
@@
For comparison, UNICORN and PROGRAPHER all perform worse than METANOIA in nearly every metric. We learned that there are two reasons why UNICORN’s performance is worse than reported in the original paper: (1) UNICORN was trained using benign datasets under DARPA TC that were not publicly available, leading to bias in the model. (2) UNICORN did not enforce that the testing graphs occur after training and validation, which is a typical Data Snooping [arp2022datasnooping]. While PROGRAPHER was not open-sourced, we replicated it based on the description of the paper. Keeping the hyper-parameters and necessary components unchanged, we believe the main reason for PROGRAPHER’s poorer performance compared to the original paper is that the encoder (Graph2Vec) cannot generate suitable embedding vectors. We primarily used Doc2Vec [le2014doc2vec] provided by Gensim [gensim] in combination with RSG mentioned in PROGRAPHER to generate embedding vectors for snapshots, but the performance was unsatisfactory. We attribute the result to our lack of expertise rather than the original authors, and we intend to open-source the reproduction code of PROGRAPHER.
@@ VI-D3 Node-level
THREATRACE is a detection system that assesses the anomalousness of nodes using a multi-model framework. While METANOIA generates a reconstruction graph of the attack scenario corresponding to the malicious windows, allowing for evaluation at the node level for both. THREATRACE [threatrace] is an open-source system, enabling us to run it directly for comparison. To maintain fairness, we adopted THREATRACE’s way of computing metrics and selected a subset of the DARPA TC dataset used by THREATRACE.
@@ VI-E RQ2: The Importance of METANOIA’s Components
Can the RN Pool be contaminated? From Table VII, we can observe that the constructed RN Pool has no intersection with the attack nodes across almost all datasets. This is attributed to METANOIA’s stringent suspicious state transfer rules in constructing the RN Pool, thereby enabling METANOIA to avoid learning malicious behaviors and introducing false negatives. Prior work [sleuth, xiong2020conan] suggests that suspicious state transfer captures the essence of intrusion attacks, and METANOIA utilizes this rule to inversely identify rehearsal nodes that can provide replay information of critical scenarios. Additionally, in the case of E3-ClearScope, due to the same reason (absence of events related to process creation), the suspicious state can’t transfer effectively, resulting in all attack-related nodes being included in the RN Pool. We attribute this to issues with the auditing system.
@@ VI-F RQ3: The Effectiveness of Reconstructing Attack Scenarios
Previous researches [kairos, alsaheel2021atlas, xu2022depcomm] indicate that reconstructing attack scenarios is an essential capability for existing detection systems. Unlike rule-based detection methods, anomaly-based detection systems focus more on this ability. The reason for this is that anomaly-based detection systems inherently have a certain amount of false positives, requiring involvement from security analysts. Reconstructing a complete and concise attack scenario can effectively assist security analysts in eliminating false positives and reducing time overhead.
@@ VII Discussion
Evasion Attack. Evasion attack refers to the actions taken by the attacker to mimic benign behavior to deceive detection systems during an attack. While evasion attacks pose a threat to all PIDSes, evading anomaly-based PIDSes is not straightforward. Therefore, to mimic benign behavior, attackers need to carefully orchestrate their attack actions, ensuring that malicious behavior exhibits benign features at the log level while maintaining their attack objectives. This requires attackers to have a deep understanding of both the target system’s benign behavior and the mechanics of running PIDSes. Previous work [SIGL] has demonstrated that existing graph-based adversarial attacks cannot evade the detection of PIDSes because the provenance graphs have more structural features and time constraints. A recent robustness study [goyal2023sometimes] has indicated that evasion attacks can be achieved by automatically modifying system logs to evade existing PIDSes and have released evasion datasets based on DARPA E3-THEIA. However, KAIROS pointed out that this dataset only includes a small portion of the attack activities described in the DARPA ground truth, rather than the complete attack traces included in the original DARPA dataset. Further investigation confirmed this, and we did not proceed with further evaluation.
Poisoning Attack. If an attacker can poison the training data to contain malicious activity for model learning, then future attacks will go undetected. To the best of our knowledge, SIGL [SIGL] is the only PIDS that evaluates its robustness against data poisoning in-depth, but only for the software installation scenario. ShadeWatcher [shadewatcher] inserts attack data from one day during training and demonstrates that it can still detect the corresponding attack. We performed a similar evaluation and also obtained good detection results. METANOIA adopts a new anomaly detection paradigm that requires replaying scenarios to combat catastrophic forgetting and is therefore more sensitive to poisoning attacks. METANOIA starts from the opposite side of the suspicious information flow to avoid the discrimination paradox and thus combat the poisoning attack.
@@ VIII Related Work
In terms of detection methods, existing PIDSes can be categorized as heuristic-based and anomaly-based. Heuristic-based PIDSes [holmes, poirot, nodoze, rapsheet, sleuth] use empirical knowledge to construct matching rules to find known attacks in the provenance graph. However, they have difficulties in detecting unknown attacks and extending scale. Anomaly-based PIDSes detect intrusions by identifying deviations from normal behavior. Besides, offline PIDSes [shadewatcher, wang2020provdetector, SIGL], due to their high operational overhead, fail to achieve efficient real-time intrusion detection. Online PIDSes [han2020unicorn, threatrace, prographer] relies on a fixed-size training set and cannot effectively address the problem of concept drift. UNICORN [han2020unicorn] combats concept drift by gradual forgetting but causes key information to decay with duration. KAIROS [kairos] relies on the intervention of analysts to re-train the model to combat concept drift, resulting in breaking attack-agnostic and high training costs. Although KAIROS acknowledges this issue in the paper, it does not provide a targeted solution, leaving it for future work.
Moreover, previous PIDSes [pasquier2017practical, FRAPpuccino, streamspot, han2020unicorn] required manual investigation of large-scale provenance graphs, rendering them impractical for real-world use. Recently, THREATRACE [threatrace] and SHADEWATCHER [shadewatcher] began identifying anomalies at the node level, but they cannot reconstruct complete attack scenarios like METANOIA. SIGL [SIGL] attempts to generate graphs for software installation, but it struggles to scale to full-system intrusion detection. PROGRAPHER [prographer], on the other hand, sorts the root subgraphs (RSGs) in snapshots by anomaly severity to help narrow down the scope of the investigation but loses the connection between multiple snapshots. Meanwhile, KAIROS [kairos] attempts to reconstruct attack scenarios in anomaly queues using community discovery algorithms but can only generate multiple star-shaped small attack scenarios (according to its open-source results).